# Particle Thermal Inertia Delays the Onset of Convection in Particulate Rayleigh–Bénard System

### v1

@@ 2 The particulate Rayleigh-Bénard model system
A few additional observations are in order. The operators $D_{t}()=\partial_{t}()+\textbf{u}\cdot\nabla()$ and $d_{t}()=\partial_{t}()+\textbf{w}\cdot\nabla()$ denote the material derivatives for the fluid and particulate phases, respectively, where $\nabla()$ is the spatial gradient operator. Three hydrodynamic forces are considered: (i) Stokes drag, (ii) added-mass correction to fluid acceleration, and (iii) buoyancy. The drag force is characterized by the viscous response time $\tau_{p}=\textrm{d}_{p}^{2}/(12\,\nu\,\beta)$, where $\nu$ is the fluid kinematic viscosity. The added-mass contribution depends on the modified density ratio $\beta=3\rho/(\rho+2\rho_{p})$, where $\rho$ is the fluid density. The particle temperature is assumed spatially uniform (lumped approximation), with relaxation toward the local fluid temperature governed by the timescale $\tau_{T}=(3-\beta)/(2\beta)\,\epsilon\,\textrm{d}_{p}^{2}/(12\kappa)$, where $\kappa$ is the fluid thermal diffusivity and $\epsilon=c_{Pp}/c_{P}$ is the ratio of particle to fluid specific heat capacity at constant pressure. The remaining constants include the fluid volumetric thermal expansion coefficient $\beta_{T}$ at the reference temperature $T_{r}$, the gravitational acceleration vector g, and the pressure field $p(\textbf{x},t)$.
@@ 2.1 Dimensionless system
where $Ra$ is the Rayleigh number, representing the ratio of buoyant to diffusive effects, $Pr$ is the Prandtl number, expressing the relative importance of momentum and thermal diffusivities, $Ga$ is the Galileo number, quantifying the balance between gravitational and viscous forces, and $\Phi$ is the particle-to-domain size ratio. Although $Ga$ does not depend on particle properties, it becomes a relevant control parameter when particle–fluid coupling is included. In dimensionless form, the terminal particle velocity is $\mathbf{W}_{T}=(1-\beta)/\beta\,(\Phi^{2}/12)\,Ga\,Pr\,\hat{\mathbf{Z}}$ . Together with the modified density ratio $\beta$, the particle inlet flux $\mathcal{J}$, and the inlet velocity and temperature $(W^{\ast},\Theta_{p}^{\ast})$, these quantities define the full set of control parameters. In total, the model is governed by nine parameters: three associated with the fluid phase $(Ra,Pr,Ga)$ and six associated with the particulate phase $(\Phi,\beta,\epsilon,\mathcal{J},W^{\ast},\Theta_{p}^{\ast})$.
@@ 3.2 Perturbation equations
Equations (22)–(28) are solved numerically using the shooting method, with the critical conditions ($Ra_{c},\lambda_{c},k_{c}$) identified from the sensitivity of $Ra$ to variations in $k$ (Alves et al., 2019). To verify the results, a matrix-forming approach is also employed. In this formulation, the differential system is discretized with a fourth-order finite-difference scheme and cast as a generalized algebraic eigenvalue problem. The resulting matrix system is then solved using the Arnoldi method with a shift-and-invert spectral transformation (Souza et al., 2021). Both methods show excellent quantitative agreement, as shown in Table 1.
@@ 4 Results and discussion
In this study, some parameters are kept constant: $Pr=5$, $Ga=9.8\times 10^{9}\$ [^1] It is worth noting that this value matches the one used in our previous works: (Raza et al., 2025), where it was inadvertently reported incorrectly, and (Raza et al., 2024), where it appeared within the parameter $\Lambda=GaPr^{2}$, not used in the current study., and $\Phi=10^{-2}$. The critical Rayleigh number $Ra_{c}$ serves as the primary control variable for the stability analysis. The parameters varied are $\beta$, $\epsilon$, $\Theta_{p}^{\ast}$, $\mathcal{J}$, and $\mathbf{W}^{\ast}$. Unless otherwise stated, the particle flux and injection velocity are set to $\mathcal{J}_{0}=533.3$ and $\mathbf{W}^{\ast}=0.5,\mathbf{W}_{T}$, respectively, whereas the particle injection temperature $\Theta_{p}^{\ast}$ equals the temperature of the wall from which the particles are introduced—hot ($\Theta_{p}^{\ast}=1$) for light particles injected from below, and cold ($\Theta_{p}^{\ast}=0$) for heavy particles injected from above.
@@ 4.1 Influence of the specific heat capacity ratio
Figure 2 shows the dependence of the critical Rayleigh number $Ra_{c}$ on the specific heat capacity ratio $\epsilon$ for representative values of $\beta$ corresponding to heavy and light particles. The parameter $\epsilon$ quantifies the efficiency of heat exchange between particles and fluid, thereby controlling the strength of thermal coupling. For small $\epsilon$, the particle heat capacity is negligible compared to that of the fluid, and the particles rapidly equilibrate with the surrounding temperature field. In contrast, for large $\epsilon$, particles act as distributed heat sources or sinks, influencing the local thermal field while maintaining their own temperature. The mechanical coupling between the two phases remains unaffected, as it is independent of $\epsilon$. This separation of mechanical and thermal effects distinguishes the present parametrization from earlier pRB studies.
@@ 4.3 Influence of the particulate mass density
In the light-particle regime ($\beta>1$), distinct trends emerge. As $\beta$ increases beyond unity, the three curves separate, showing that the influence of the specific heat capacity ratio $\epsilon$ becomes significant. Consistent with previous results, larger $\epsilon$ values, i.e. stronger particle heat capacity relative to the fluid—lead to higher critical Rayleigh and wavenumbers, indicating increased stability and smaller dominant structures, respectively. When $\beta$ is sufficiently large, however, all curves shift downward. For instance, when $\epsilon=0.1$, $Ra_{c}$ falls below the single-phase Rayleigh–Bénard threshold. Varying $\beta$ therefore modifies both mechanical and thermal coupling effects. This dual influence can be seen directly in Equation (6), where the limit $\beta\to 3$ is equivalent to $\epsilon\to 0$. For $\beta=3$, conventionally referred to as the “bubble” case, the thermal coupling vanishes.

### v2

@@ 2 The particulate Rayleigh-Bénard model system
A few additional observations are in order. The operators $D_{t}()=\partial_{t}()+\textbf{u}\cdot\nabla()$ and $d_{t}()=\partial_{t}()+\textbf{w}\cdot\nabla()$ denote the material derivatives for the fluid and particulate phases, respectively, where $\nabla()$ is the spatial gradient operator. Three hydrodynamic forces are considered: (i) Stokes drag, (ii) added-mass correction to fluid acceleration, and (iii) buoyancy. The drag force is characterized by the viscous response time $\tau_{p}=\textrm{d}_{p}^{2}/(12\,\nu\,\beta)$, where $\nu$ is the fluid kinematic viscosity. The added-mass contribution depends on the modified density ratio $\beta=3\rho/(\rho+2\rho_{p})$, where $\rho$ is the fluid density. The particle temperature is assumed spatially uniform (lumped approximation), with relaxation toward the local fluid temperature governed by the timescale $\tau_{T}=(3-\beta)/(2\beta)\,\epsilon\,\textrm{d}_{p}^{2}/(12\kappa)$, where $\kappa$ is the fluid thermal diffusivity and $\epsilon=c_{Pp}/c_{p}$ is the ratio of particle to fluid specific heat capacity at constant pressure. The remaining constants include the fluid volumetric thermal expansion coefficient $\beta_{T}$ at the reference temperature $T_{r}$, the gravitational acceleration vector g, and the pressure field $p(\textbf{x},t)$. We stress that the model takes into account thermal and mechanical couplings between the fluid and the particles (two-way) but neglects particle-particle (four-way) couplings, e.g. contact collisions and lubrication forces.
@@ 2.1 Dimensionless system
where $Ra$ is the Rayleigh number, representing the ratio of buoyant to diffusive effects, $Pr$ is the Prandtl number, expressing the relative importance of momentum and thermal diffusivities, $Ga$ is the Galileo number, quantifying the balance between gravitational and viscous forces, and $\Phi$ is the particle-to-domain size ratio. Although $Ga$ does not depend on particle properties, it becomes a relevant control parameter when particle–fluid coupling is included. In dimensionless form, the terminal particle velocity is $\mathbf{W}_{T}=(\beta-1)/\beta\,(\Phi^{2}/12)\,Ga\,Pr\,\hat{\mathbf{Z}}$ . Together with the modified density ratio $\beta$, the particle inlet flux $\mathcal{J}$, and the inlet velocity and temperature $(W^{\ast},\Theta_{p}^{\ast})$, these quantities define the full set of control parameters. In total, the model is governed by nine parameters: three associated with the fluid phase $(Ra,Pr,Ga)$ and six associated with the particulate phase $(\Phi,\beta,\epsilon,\mathcal{J},W^{\ast},\Theta_{p}^{\ast})$.
@@ 3.2 Perturbation equations
Equations (22)–(28) are solved numerically using the shooting method, with the critical conditions ($Ra_{c},\lambda_{c},k_{c}$) identified by minimizing $Ra$ with respect to $k$ (Alves et al., 2019). To verify the results, a matrix-forming approach is also employed. In this formulation, the differential system is discretized with a fourth-order finite-difference scheme and cast as a generalized algebraic eigenvalue problem. The resulting matrix system is then solved using the Arnoldi method with a shift-and-invert spectral transformation (Souza et al., 2021). Both methods show excellent quantitative agreement, as shown in Table 1.
@@ 4 Results and discussion
In this study, some parameters are kept constant: $Pr=5$, $Ga=1.92\times 10^{10}\$ [^1] It is worth noting that this value matches the one used in our previous works: (Raza et al., 2025), where it was inadvertently reported incorrectly, and (Raza et al., 2024), where it appeared within the parameter $\Lambda=GaPr^{2}$, not used in the current study., and $\Phi=10^{-2}$. The critical Rayleigh number $Ra_{c}$ serves as the primary control variable for the stability analysis. The parameters varied are $\beta$, $\epsilon$, $\Theta_{p}^{\ast}$, $\mathcal{J}$, and $\mathbf{W}^{\ast}$. Unless otherwise stated, the particle flux and injection velocity are set to $\mathcal{J}_{0}=533.3$ and $\mathbf{W}^{\ast}=0.5\mathbf{W}_{T}$, respectively, whereas the particle injection temperature $\Theta_{p}^{\ast}$ equals the temperature of the wall from which the particles are introduced—hot ($\Theta_{p}^{\ast}=1$) for light particles injected from below, and cold ($\Theta_{p}^{\ast}=0$) for heavy particles injected from above.
@@ 4.1 Influence of the specific heat capacity ratio
Figure 2 shows the dependence of the critical Rayleigh number $Ra_{c}$ on the specific heat capacity ratio $\epsilon$ for representative values of $\beta$ corresponding to heavy and light particles. The parameter $\epsilon$ quantifies the efficiency of heat exchange between particles and fluid, thereby controlling the strength of thermal coupling. For small $\epsilon$, the particle heat capacity is negligible compared to that of the fluid, and the particles rapidly equilibrate with the surrounding temperature field. In contrast, for large $\epsilon$, particles act as distributed heat sources or sinks, influencing the local thermal field while maintaining their own temperature. While the limit $\epsilon\rightarrow\infty$ is idealized, large but finite values of $\epsilon$ can be physically relevant. For example, the specific heat capacity ratio of ice particles in air is $\sim 2$ while cryogenic liquid hydrogen drops in mercury (Hg) it is $\sim 100$. In contrast, for solid particles such as sand in water ratio becomes $\sim 0.19$ and for metal particles such as cooper (Cu) in water the ratio is $\sim 0.09$. The mechanical coupling between the two phases remains unaffected, as it is independent of $\epsilon$. This separation of mechanical and thermal effects distinguishes the present parametrization from earlier pRB studies.
@@ 4.3 Influence of the particulate mass density
In the light-particle regime ($\beta>1$), distinct trends emerge. As $\beta$ increases beyond unity, the three curves separate, showing that the influence of the specific heat capacity ratio $\epsilon$ becomes significant. Consistent with previous results, larger $\epsilon$ values, i.e. stronger particle heat capacity relative to the fluid—lead to higher critical Rayleigh and wavenumbers, indicating increased stability and smaller dominant structures, respectively. When $\beta$ is sufficiently large, however, all curves shift downward. For instance, when $\epsilon=0.1$, $Ra_{c}$ falls below the single-phase Rayleigh–Bénard threshold. Varying $\beta$ therefore modifies both mechanical and thermal coupling effects. This dual influence can be seen directly in Equation (6), where the limit $\beta\to 3$ is equivalent to $\epsilon\to 0$. In this limit, the particles instantaneously adjust their temperature to that of the surrounding fluid. The resulting dimensional equation governing the fluid temperature is
$D_{t}T=\frac{\kappa}{1-\alpha},\nabla^{2}T.$ (30)
This result indicates that bubbles ($\beta=3$) and particles with vanishing specific heat capacity ($\epsilon=0$) modify the thermal properties of the fluid solely through an excluded-volume effect. In particular, the effective thermal diffusivity of the fluid is enhanced by a factor $1/(1-\alpha)$. In the present study, this correction is negligible, since the mean particle concentration is $O(10^{-3})$.
@@ 5 Conclusion
Funding. This work was partially supported by the CAPES-COFECUB programme (project number: Ph1097/26), funded by the French Ministry for Europe and Foreign Affairs, the French Ministry for Higher Education and CAPES.